OpenStudy (anonymous):

Use the Law of Sines and Cosines to solve the following triangle? A = 10.3 degrees, C = 143.7 degrees, c = 48.3

myininaya (myininaya):



myininaya (myininaya):

\[B=180-(A+C)\]

myininaya (myininaya):

\[\frac{\sin(A)}{a}=\frac{\sin(C)}{c}\]

myininaya (myininaya):

\[b^2=a^2+c^2-2 a c \cos(B)=> b=\sqrt{a^2+c^2- 2 a c \cos(B)}\]